# Phenomenology of laminar acoustic streaming jets

### v1

@@ IV.3 Viscosity-dominated flow region downstream the velocity peak
In Ref. [31] a similar longitudinal rate of change is assumed for $R_{jet}$ and $R_{beam}$, so the variations of $U_{x,d}$ could only arise from the attenuation of the beam. This is not observed here (Fig. 3 (b)), and beam attenuation is too weak to be sole responsible for the decay of $u_{x}$ downstream $x_{p}$ (Fig. 5).
Figure 8: On-axis velocity profiles downstream the velocity peak $x_{p}$, obtained for different $Gr_{ac}$. The quantity plotted on the $x$-axis includes the longitudinal variations of the acoustic force caused by both diffraction and attenuation of the beam (see the right-hand side of Eq. (18)). The order of magnitude $U_{x,d}$ of the on-axis velocity downstream $x_{p}$ (Eq. (18)), is shown in black. The colored solid lines correspond to the numerical data. The dashed lines represent the self-similar jet solution $U_{x,S}$ of Schlichting [51] given by Eq. (20) and based on the maximum axial momentum flux $M=\pi LGr_{ac}/\left(32N\right)=\pi L_{\alpha}Gr_{ac}/16$ asymptotically acquired by the jet. The red vertical line locates the force attenuation distance $L_{\alpha}=L/\left(2N\right)$.
It is thus of primary importance to rely on a good estimate of the axial jet profile to assess viscosity effects. Contrarily to Ref. [31], we do not assume $R_{jet}=R_{beam}$, but instead base $R_{jet}$ on equation (13) describing the build up of the jet momentum flux along $x$. Equation (17) then gives:
@@
As the considered flow region is located downstream $L_{F}$, the enlargement of the beam due to diffraction is significant and needs to be taken into account in the definition of $X$. This means that $R_{beam}$ can no longer be considered as constant, but shall instead follow the linear expansion given by Eq. (6). Thus, from Eq. (18), the longitudinal variations of $u_{x}$ are caused by both diffraction and attenuation. As shown in Fig. 8, Eq. (18) (plotted as a solid black line) accurately describes the longitudinal decrease of $u_{x}$ observed in the numerical simulations up to $x\approx L_{\alpha}$.
@@ IV.4 Negligible acoustic force region far downstream the velocity peak
At greater $x$-coordinates, the acoustic force amplitude keeps decreasing due to attenuation. Its magnitude relative to inertia keeps reducing with $x$, to the point where both become comparable for $x\simeq L_{\alpha}$.
Another approach is thus required for $x\gg L_{\alpha}$, a region where Eq. (18) is no more valid. In this fluid region, most of the energy of the acoustic beam has been attenuated, meaning that the forcing is negligible. The jet thus obeys a balance between inertial and viscous forces for $x\gg L_{\alpha}$. Furthermore, as shown in Sec. III, the beam attenuation causes the axial momentum of the acoustic streaming jet to asymptotically reach the constant value $M=\pi LGr_{ac}/\left(32N\right)=\pi L_{\alpha}Gr_{ac}/16$. The situation thus becomes similar to the case of a straight jet of constant axial momentum flux $M$ emerging from an orifice, described by the boundary layer equation:
@@
$U_{x,S}=\frac{3L_{\alpha}Gr_{ac}}{128x}\,.$ (20)
Equation (20), plotted as dashed lines in Fig. 8, captures the asymptotic behavior of the on-axis velocity at $x\gg L_{\alpha}$: for $Gr_{ac}=2\times 10^{3}$ for instance, $U_{x,S}$ and $u_{x}$ differ by 0.6 % just upstream the impingement.
@@ V Conclusion
We presented a numerical study of an Eckart-type acoustic streaming jet. The flow develops in a wide and elongated cylindrical cavity entirely filled with a viscous liquid. The acoustic source is a plane circular transducer, of axis aligned with the cylinder axis, placed at one end of the cavity in direct contact with the fluid. It radiates an axisymmetric beam-shaped acoustic field yielding a body force that drives the flow. An important aspect of this work is that the characteristic attenuation length $L_{\alpha}$ of the acoustic force is smaller but comparable to the cavity length, so that the jet features both a region close to the source with prominent acoustic force, and regions further away where it is not. The diameter of the cavity is sufficiently large to avoid the influence of the lateral wall on the jet. The range of acoustic force magnitudes, measured by the acoustic Grashof number $Gr_{ac}$, is limited to moderate values, so that the resulting streaming jet remains axisymmetric, steady and laminar. Hence, this configuration provides a generic representation of applications where a transducer is directed into a large volume so as to act over a distance greater than the attenuation length but smaller than the size of the cavity (this would prove particularly useful in metallurgical and crystal growth applications [52, 19]). In this type of configuration, acoustic attenuation plays a central role for the modeling of the acoustic field even at small distance from the source. Indeed, the error incurred on the injected momentum flux when neglecting attenuation is already 10 % at a distance $0.2L_{\alpha}$ from the source, and further increases with the on-axis distance $x$ from the transducer. Despite being frequently used in the literature, the unattenuated beam approximation leads to significant errors: it severely overestimates the jet velocity along its axis, which in turn affects the flow structure by reducing the impact of viscous diffusion on the radial spreading of the jet.
@@ A.1 Governing equations of acoustics
Although acoustic streaming refers to an incompressible flow, it relies on the attenuation of sound waves, which are a compressible process [38]. Based on the assumptions listed in Sec. II.2, the acoustic propagation problem can be described by the compressible Navier-Stokes and continuity equations linearized around a resting fluid of density $\rho$, together with an isentropic equation of state [36, 37]:
@@
where $p_{ac}$ and $\rho_{ac}$ are the acoustic fluctuations of pressure and density, respectively, $c$ is the speed of sound and ${\tau=\frac{1}{\rho c^{2}}\left(\frac{4}{3}\mu+\eta\right)}$.
We shall also use the wave equation for $\bm{u_{ac}}$:
$\bm{\nabla}^{2}\bm{u_{ac}}+k^{2}\bm{u_{ac}}=\bm{0}\,,$ (22)
which is obtained by combining the divergence of Eq. (21a) with the time-derivative of Eq. (21b). In Eq. (22), the effects of attenuation appear through the imaginary part of the complex wavenumber $k$ defined as:
@@
Finally, under the assumptions (ii) and (iii) of Sec. II.2), the momentum equation (21a) can be rewritten to give:
$j\frac{\rho c^{2}k^{2}}{\omega}\bm{u_{ac}}=-\nabla p_{ac}\,,$ (26)
@@
where * is the complex conjugate.
@@ A.2 Integral relations for the acoustic intensity field radiated by a plane circular transducer
As for the unattenuated case, the vibrating surface $S_{t}$ of the transducer is modeled as a plane disc made of an infinite number of point sources [36, 37], each centered on an elementary surface $dS$ and creating a flow rate $Q_{0}$ through each $dS$ (Fig. 10).
@@
Figure 10: Sketch of the circular transducer of diameter $D_{s}$ (gray), modeled as a plane disc made of an infinite number of point sources. Each source is centered on an elementary surface $dS$ and radiates spherical acoustic waves contributing to the net acoustic pressure and velocity (hence intensity) levels at $P$ in the fluid domain.
@@
On $S_{t}$, $Q_{0}=\bm{u_{0}}\cdot\bm{n_{t}}dS$, with $\bm{u_{0}}$ the vibrating velocity of the transducer and $\bm{n_{t}}$ the normal unit vector. Because the superposition principle applies, $p_{ac}$ in the fluid domain is a linear combination of all the pressure waves emitted by each point source and is thus obtained by integrating Eq. (29) over $S_{t}$:
[…]
in which $Q_{0}$ has been doubled, since each point source radiates waves in the $x>0$ semi-infinite domain only. As the transducer vibrates uniformly, the acoustic fields are axisymmetric and thus need to be calculated in a single constant-$\theta$ half-plane only. For simplicity, we shall consider $\theta=0$, so that the vectorial distance $\bm{d}$ between a point source at $\left(r^{\prime},\theta^{\prime}\right)$ on $S_{t}$ and the point at $\left(x,r\right)$ where $p_{ac}$ is evaluated (Fig. 10) reduces to:
$\bm{d}=\left(x\,\bm{e_{x}}+\left[r-r^{\prime}\cos\left(\theta^{\prime}\right)\right]\,\bm{e_{r}}-r^{\prime}\sin\left(\theta^{\prime}\right)\bm{e_{\theta}}\right)\,.$
@@
[…]
in which $Q_{0}$ has again been doubled for the same reason as in Eq. (30). From Eq. (27), it appears that the temporal variations of $p_{ac}$ and $\bm{u_{ac}}$ do not affect $\bm{I}$; we can thus drop $e^{j\omega t}$ from Eqs. (30) and (31).
@@ A.3 Analytical expressions of the acoustic fields on the transducer axis
Except at particular points in the fluid domain, the integral expressions for $p_{ac}$ and $\bm{u_{ac}}$ derived in the previous section cannot be computed analytically in the entire $x>0$ domain, and shall thus be evaluated numerically. It is nevertheless possible to obtain analytical expressions for $p_{ac}$, $\bm{u_{ac}}$ and $\bm{I}$ on the acoustic axis, which prove useful when checking the numerical computations of Eqs. (30) and (31).
We shall first evaluate $p_{ac}$ at $r=0$: there, $\|\bm{d}\|=x^{2}+r^{\prime 2}$, so that Eq. (30) can be calculated analytically to give:
@@
Contrarily to $p_{ac}$, Eq. (31) cannot be calculated analytically even for $r=0$. However, ${\bm{u_{ac}}(x,r=0)=u_{ac,x}(x,r=0)\bm{e_{x}}}$ can be computed using Eq. (26), in which $\nabla p_{ac}$ is calculated from Eq. (32). The resulting expression of $u_{ac,x}$ can finally be used with Eq. (27) to get the on-axis acoustic intensity $I_{x}(x,r=0)$:

### v2

@@ IV.3 Viscosity-dominated flow region downstream the velocity peak
In Ref. [31] a similar longitudinal rate of change is assumed for $R_{jet}$ and $R_{beam}$, so the variations of $U_{x,d}$ could only arise from the attenuation of the beam. This is not observed here (Fig. 3 (b)), and beam attenuation is too weak to be the sole reason for the decay of $u_{x}$ downstream $x_{p}$ (Fig. 5).
Figure 8: On-axis velocity profiles downstream the velocity peaks $x_{p}$, which are represented by the black points. THe profiles are shown for different $Gr_{ac}$. The quantity plotted on the $x$-axis includes the longitudinal variations of the acoustic force caused by both diffraction and attenuation of the beam (see the right-hand side of Eq. (18)). The red vertical line locates the force attenuation distance $L_{\alpha}=L/\left(2N\right)$. The scaling law (18) for the on-axis velocity $U_{x,d}$ at $x_{p}\leq x\lessapprox L_{\alpha}$ is represented by the thick straight black line. The colored solid lines correspond to the numerical data. The dashed curves represent the self-similar jet solution $U_{x,S}$ of Schlichting [53] given by Eq. (20) and based on the maximum axial momentum flux $M=\pi LGr_{ac}/\left(32N\right)=\pi L_{\alpha}Gr_{ac}/16$ asymptotically acquired by the jet.
It is thus of primary importance to rely on a good estimate of the axial jet profile to assess viscosity effects. Contrary to Ref. [31], we do not assume $R_{jet}=R_{beam}$, but instead base $R_{jet}$ on equation (13) describing the build up of the jet momentum flux along $x$. Equation (17) then gives:
@@
As the considered flow region is located downstream $L_{F}$, the enlargement of the beam due to diffraction is significant and needs to be taken into account in the definition of $X$. This means that $R_{beam}$ can no longer be considered as constant, but shall instead follow the linear expansion given by Eq. (6). Thus, from Eq. (18), the longitudinal variations of $u_{x}$ are caused by both diffraction and attenuation.
As shown in Fig. 8, $U_{x,d}$ (plotted as a thick straight black line) correctly captures the slope the on-axis $u_{x}$ (solid colored curves) on $x_{p}\leq x\leq L_{\alpha}$ and for all $Gr_{ac}$. Besides, $U_{x,d}$ differs from $u_{x}$ by a factor at most 3, meaning that Eq. (18) also correctly estimates the order of magnitude of $u_{x}$ on that $x$ interval.
@@ IV.4 Negligible acoustic force region far downstream the velocity peak
Beyond $L_{\alpha}$, the magnitude of the acoustic force relative to inertia decreases due to attenuation, causing $u_{x}$ to deviate from $U_{x,d}$. Equation (18) is thus unable to describe the on-axis $u_{x}$ over the entire $x\geq x_{p}$ jet region, and another approach is needed to recover the on-axis profile of $u_{x}$ at $x\gg L_{\alpha}$.
At $x\gg L_{\alpha}$, most of the energy of the acoustic beam has been attenuated, meaning that the forcing is negligible whatever the $Gr_{ac}$. As the pressure forces also are negligible, the jet necessarily satisfies a balance between inertial and viscous forces at $x\gg L_{\alpha}$ according to the Navier-Stokes equations (3b). Furthermore, as shown in Sec. III, the beam attenuation causes the axial momentum of the acoustic streaming jet to asymptotically reach the constant value $M=\pi LGr_{ac}/\left(32N\right)=\pi L_{\alpha}Gr_{ac}/16$. The situation thus becomes similar to the case of a straight jet of constant axial momentum flux $M$ emerging from an orifice and described by the boundary layer equation:
@@
$U_{x,S}=\frac{3L_{\alpha}Gr_{ac}}{128x}\,$, (20)
which is thus expected to accurately describe the on-axis jet velocity as $x\rightarrow+\infty$.
The scaling law (20) for $U_{x,S}$ is plotted as dashed curves for all $Gr_{ac}$ in Fig. 8 alongside the numerical velocity profiles (solid curves). As expected, the decrease of the acoustic forcing relative to inertia as $x$ increases (or equivalently, as $X$ reduces) causes the numerical profiles to come closer to $U_{x,S}$. This strong agreement confirms the suitability of $U_{x,S}$ for describing the on-axis jet velocity as $x\rightarrow+\infty$.
@@ V Conclusion
We presented a numerical study of an Eckart-type acoustic streaming jet representative of former experiments [27, 29, 38, 31]. The flow develops in a wide and elongated cylindrical cavity entirely filled with a viscous liquid. The acoustic source is a plane circular transducer, of axis aligned with the cylinder axis, placed at one end of the cavity in direct contact with the fluid. It radiates an axisymmetric beam-shaped acoustic field yielding a body force that drives the flow. An important aspect of this work is that the characteristic attenuation length $L_{\alpha}$ of the acoustic force is smaller but comparable to the cavity length, so that the jet features both a region close to the source with prominent acoustic force, and regions further away where it is not. The diameter of the cavity is sufficiently large to avoid the influence of the lateral wall on the jet. The range of acoustic force magnitudes, measured by the acoustic Grashof number $Gr_{ac}$, is limited to moderate values, so that the resulting streaming jet remains axisymmetric, steady and laminar. Hence, this configuration provides a generic representation of applications where a transducer is directed into a large volume so as to act over a distance greater than the attenuation length but smaller than the size of the cavity (this would prove particularly useful in metallurgical and crystal growth applications [54, 19]). In this type of configuration, acoustic attenuation plays a central role for the modeling of the acoustic field even at small distance from the source. Indeed, the error incurred on the injected momentum flux when neglecting attenuation is already 10 % at a distance $0.2L_{\alpha}$ from the source, and further increases with the on-axis distance $x$ from the transducer. Despite being frequently used in the literature, the unattenuated beam approximation leads to significant errors: it severely overestimates the jet velocity along its axis, which in turn affects the flow structure by reducing the impact of viscous diffusion on the radial spreading of the jet.
@@ A.1 Governing equations of acoustics
Although acoustic streaming refers to an incompressible flow, it relies on the attenuation of sound waves, which are a compressible process [39]. Based on the assumptions listed in Sec. II.2, the propagation of linear sound waves far from solid boundaries can be described by the irrotational compressible Navier-Stokes and continuity equations linearized around a resting fluid of density $\rho$, together with an isentropic equation of state [36, 37]:
@@
where $p_{ac}$ and $\rho_{ac}$ are the acoustic fluctuations of pressure and density, respectively, $c$ is the speed of sound and ${\tau=\frac{1}{\rho c^{2}}\left(\frac{4}{3}\mu+\eta\right)}$. The propagation of linear attenuated sound waves is then governed by a wave equation for $\bm{u_{ac}}$:
$\frac{1}{c^{2}}\frac{\partial^{2}\bm{u_{ac}}}{\partial t^{2}}-\bm{\nabla}^{2}\bm{u_{ac}}=\tau\frac{\partial}{\partial t}\left(\bm{\nabla}^{2}\bm{u_{ac}}\right)\,$,
which is obtained by (i) taking the time-derivative of Eq. (21a), (ii) taking the gradient of Eq. (21b), (iii) combining the obtained equations and (iv) making an extensive use of $\bm{\nabla}\times\bm{u}_{ac}=\bm{0}$. For a monochromatic wave $\bm{u_{ac}}=\bm{\widehat{u}_{ac}}e^{j\omega t}$, the above wave equation then reduces to a complex Helmholtz equation:
$\bm{\nabla}^{2}\bm{\widehat{u}_{ac}}+k^{2}\bm{\widehat{u}_{ac}}=\bm{0}\,.$ (22)
In Eq. (22), the effects of attenuation appear through the imaginary part of the complex wavenumber $k$ defined as:
@@
Finally, under the assumptions (ii) and (iii) of Sec. II.2, the momentum equation (21a) can be rewritten to give:
$j\frac{\rho c^{2}k^{2}}{\omega}\bm{\widehat{u}_{ac}}=-\nabla\widehat{p}_{ac}\,$, (26)
@@
where is the complex conjugate.
@@ A.2 Integral relations for the acoustic intensity field radiated by a plane circular transducer
To obtain the integral expression for $p_{ac}$ and $\bm{u_{ac}}$ (and hence to infer $\bm{I}$), we shall closely follow the derivation of Blackstock [37] and adapt it to account for sound attenuation. The idea is to model the vibrating surface $S_{t}$ of the transducer as a plane disc made of an infinite number of point sources. Each of these sources is centered on an elementary surface $\mathrm{d}S$ (Fig. 10) and drives a net flow rate $Q_{0}$ through that surface. We shall thus first focus on the spherical pressure wave emitted by a single point source, and determine how the wave amplitude relates to $Q_{0}$. The acoustic pressure and velocity levels in the fluid domain are then obtained by summing up the spherical waves radiated by all point sources on $S_{t}$. We checked that the same pressure Rayleigh integral as Eq. (30) can be obtained using Green’s functions (see Ref. [63] for the derivation in the context of unattenuated acoustics).
@@
Figure 10: Sketch of the circular transducer of diameter $D_{s}$ (gray), modeled as a plane disc made of an infinite number of point sources. Each source is centered on an elementary surface $\mathrm{d}S$ and radiates spherical acoustic waves contributing to the net acoustic pressure and velocity (hence intensity) levels at $P$ in the fluid domain.
@@
On $S_{t}$, $Q_{0}=\bm{u_{0}}\cdot\bm{n_{t}}\mathrm{d}S$, with $\bm{u_{0}}$ the vibrating velocity of the transducer and $\bm{n_{t}}$ the normal unit vector. Because the superposition principle applies, $p_{ac}$ in the fluid domain is a linear combination of all the pressure waves emitted by each point source and is thus obtained by integrating Eq. (29) over $S_{t}$:
$p_{ac}=j\frac{\rho c^{2}k^{2}u_{0}}{2\pi\omega}\iint_{S_{t}}\frac{e^{j\left(\omega t-k\|\bm{d}\|\right)}}{\|\bm{d}\|}\mathrm{d}S\,$, (30)
in which $Q_{0}$ has been doubled, since each point source radiates waves in the $x>0$ semi-infinite domain only [37]. Note that the same result can be more rigorously obtained using the formalism of Green’s functions: the factor 2 then results from the method of images applied at the transducer surface at $x=0$ (see the full derivation in Morse and Feshbach [64] for unattenuated acoustics). As the transducer vibrates uniformly, the acoustic fields are axisymmetric and thus need to be calculated in a single constant-$\theta$ half-plane only. For simplicity, we shall consider $\theta=0$, so that the vectorial distance $\bm{d}$ between a point source at $\left(r^{\prime},\theta^{\prime}\right)$ on $S_{t}$ and the point at $\left(x,r\right)$ where $p_{ac}$ is evaluated (Fig. 10) reduces to:
$\bm{d}=x\,\bm{e_{x}}+\left[r-r^{\prime}\cos\left(\theta^{\prime}\right)\right]\,\bm{e_{r}}-r^{\prime}\sin\left(\theta^{\prime}\right)\bm{e_{\theta}}\,.$
@@
$\bm{u_{ac}}=\frac{u_{0}}{2\pi}\iint_{S_{t}}\frac{\left(1+jk\|\bm{d}\|\right)}{\|\bm{d}\|^{2}}\frac{\bm{d}}{\|\bm{d}\|}e^{j\left(\omega t-k\|\bm{d}\|\right)}\mathrm{d}S\,$, (31)
in which $Q_{0}$ has again been doubled for the same reason as in Eq. (30). We stress that our formalism easily yields integral relations for both $p_{ac}$ and $\bm{u_{ac}}$, making the numerical implementation quite efficient and independent of the grid used for flow computations. Green’s function formalism on the other hand, cannot give a straightforward derivation of $\bm{u_{ac}}$ as in Eq (31). From Eq. (27), it appears that the temporal variations of $p_{ac}$ and $\bm{u_{ac}}$ do not affect $\bm{I}$; we can thus drop $e^{j\omega t}$ from Eqs. (30) and (31).
@@ A.3 Analytical expressions of the acoustic fields on the transducer axis
Except at particular points in the fluid domain, the integral expressions for $p_{ac}$ and $\bm{u_{ac}}$ derived in the previous section cannot be computed analytically in the entire $x>0$ domain, and shall thus be evaluated numerically. It is nevertheless possible to obtain analytical expressions for $p_{ac}$, $\bm{u_{ac}}$ and $\bm{I}$ on the acoustic axis; these expressions prove useful for checking the numerical computations of Eqs. (30) and (31).
We shall first evaluate $p_{ac}$ at $r=0$, where the attenuated Rayleigh integral (30) is:
$p_{ac}(x,r=0)=j\frac{\rho c^{2}k^{2}u_{0}}{2\pi\omega}e^{j\omega t}\int_{\theta^{\prime}=0}^{2\pi}\int_{r^{\prime}=0}^{D_{s}/2}\frac{e^{-jk\sqrt{x^{2}+r^{\prime 2}}}}{\sqrt{x^{2}+r^{\prime 2}}}\,r^{\prime}\,\mathrm{d}r^{\prime}\,\mathrm{d}\theta^{\prime}\,.$
Setting $X=\sqrt{x^{2}+r^{\prime 2}}$, the above integral can be computed analytically to give:
$\begin{split}p_{ac}(x,r=0)&=\frac{\rho c^{2}ku_{0}}{\omega}e^{j\omega t}\left(e^{-jkx}-e^{-jk\sqrt{x^{2}+\frac{D_{s}^{2}}{4}}}\right)\\ &=\frac{\rho c^{2}ku_{0}}{\omega}e^{j\omega t}e^{-j\frac{k}{2}\left(\sqrt{x^{2}+\frac{D_{s}^{2}}{4}}+x\right)}\left[e^{j\frac{k}{2}\left(\sqrt{x^{2}+\frac{D_{s}^{2}}{4}}-x\right)}-e^{-j\frac{k}{2}\left(\sqrt{x^{2}+\frac{D_{s}^{2}}{4}}-x\right)}\right]\\ &=j\frac{2\rho c^{2}ku_{0}}{\omega}e^{j\omega t}e^{-j\frac{k}{2}\left(\sqrt{x^{2}+\frac{D_{s}^{2}}{4}}+x\right)}\,\sin\left[\frac{k}{2}\left(\sqrt{x^{2}+\frac{D_{s}^{2}}{4}}-x\right)\right]\,.\end{split}$
Then, substituting the first equality of Eq. (23) for $k$ in the above equation finally gives:
@@
Equation (32) is an original result accounting for attenuation. To assess its validity, we checked that taking the modulus of Eq. (32) with $\alpha=0$ gives the same unattenuated $|p_{ac}|$ as in Ref. [36].
Contrary to $p_{ac}$, Eq. (31) cannot be calculated analytically even for $r=0$. However, $\bm{u_{ac}}(x,r=0)=u_{ac,x}(x,r=0)\bm{e_{x}}$ can be computed using Eq. (26), in which $\nabla p_{ac}$ is calculated from Eq. (32). The resulting expression of $u_{ac,x}$ can finally be used with Eq. (27) to get the on-axis acoustic intensity $I_{x}(x,r=0)$: